# revtex4-2Repair the float

### v1

@@ V Discussion & outlook
CER uses twirling to guarantee that the average-error channel studied has the required stochastic Pauli form. Coherent errors are averaged out as part of this process; while this leads to considerable simplification and increased efficiency in our error characterization, it can complicate the identification of certain physical error sources. Promisingly, as demonstrated here, identifying coherence in error models is still possible with CER, despite not using tailored approaches for estimating coherent and incoherent contributions to the error profile such as Ref. [45]. Our demonstrated ability to identify coherent terms without additional sequences suggesting that there may be ways to further reduce the experimental resources for assessing coherence, or highlight in advance the likely candidates for coherent noise that additional experimental sequences can target.
@@
We have demonstrated a physical-level characterization and used this data for predictions of logical performance. It is also possible to perform CER directly at the logical level. The studied hard cycles would then be logical gadgets, which may include mid-circuit measurements. There are now a number of different benchmarking routines for mid-circuit measurements  [19, 20, 21, 22] that could enable logical characterization in tandem with ongoing error correction. Such methods focus on logical noise at the expense of explicit physical-level characterizations, which has been a core focus of this work in addition to the specific application of estimating logical error rates. One challenge for physical-level characterizations as system sizes scale up is that the weight of relevant error terms increase in tandem with the system size. By working with logical code states, and taking marginalizations specifically tailored for the chosen code, it is probable that GRF models could be employed in a similar fashion, using physical-level characterizations to make logical-level predictions in larger codes scalably.
@@ Acknowledgements.
The authors would like to thank Arnaud Carignan-Dugas for helpful discussions regarding cycle error reconstruction and Evan Hockings for fruitful discussions on estimating marginal probabilities. The authors are thankful to Markus Müller for fruitful initial discussions and the Institute of Quantum Information at RWTH Aachen for their hospitality during a research visit by N.F. where this project was initiated. We gratefully acknowledge support by the European Union’s Horizon Europe research and innovation program under Grant Agreement Number 101114305 (“MILLENION-SGA1” EU Project), the US Army Research Office through Grant Number W911NF-21-1-0007, the Austrian Science Fund (FWF Grant-DOI 10.55776/F71) (SFB BeyondC), the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), under the Entangled Logical Qubits program through Cooperative Agreement Numbers W911NF-23-2-0216 and W911NF-23-2-0223. The Innsbruck team further received support from the IQI GmbH. M.R. acknowledges support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foun- dation) under Germany’s Excellence Strategy Cluster of Excellence Matter and Light for Quantum Computing (ML4Q) EXC 2004/1 390534769.This research is also part of the Munich Quantum Valley (K-8), which is supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of IARPA, the Army Research Office, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. Views and opinions expressed are those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them.
@@
Data and analysis scripts can be made available on request. Please contact robert.freund@uibk.ac.at.

### v2

@@ V Discussion & outlook
CER uses twirling to guarantee that the average-error channel studied has the required stochastic Pauli form. Coherent errors are averaged out as part of this process; while this leads to considerable simplification and increased efficiency in our error characterization, it can complicate the identification of certain physical error sources. Promisingly, as demonstrated here, identifying coherence in error models is still possible with CER, despite not using tailored approaches for estimating coherent and incoherent contributions to the error profile such as in Ref. [45]. Our demonstrated ability to identify coherent terms without additional sequences suggests that there may be ways to further reduce the experimental resources for assessing coherence, or highlight in advance the likely candidates for coherent noise that additional experimental sequences can target.
@@
We have demonstrated a physical-level characterization and used this data for predictions of logical performance. It is also possible to perform CER directly at the logical level. The studied hard cycles would then be logical gadgets, which may include mid-circuit measurements. There are now a number of different benchmarking routines for mid-circuit measurements [19, 20, 21, 22] that could enable logical characterization in tandem with ongoing error correction. Such methods focus on logical noise at the expense of explicit physical-level characterizations, which has been a core focus of this work in addition to the specific application of estimating logical error rates. One challenge for physical-level characterizations as system sizes scale up is that the weight of relevant error terms increases in tandem with the system size. By working with logical code states, and taking marginalizations specifically tailored for the chosen code, it is probable that GRF models could be employed in a similar fashion, using physical-level characterizations to make logical-level predictions in larger codes scalably.
@@ Acknowledgements.
The authors would like to thank Arnaud Carignan-Dugas for helpful discussions regarding cycle error reconstruction and Evan Hockings for fruitful discussions on estimating marginal probabilities. The authors are thankful to Markus Müller for fruitful initial discussions and the Institute for Quantum Information at RWTH Aachen for their hospitality during a research visit by N.F. where this project was initiated. We gratefully acknowledge support by the European Union’s Horizon Europe research and innovation program under Grant Agreement Number 101114305 (“MILLENION-SGA1” EU Project), the US Army Research Office through Grant Number W911NF-21-1-0007, the Austrian Science Fund (FWF Grant-DOI 10.55776/F71) (SFB BeyondC), the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), under the Entangled Logical Qubits program through Cooperative Agreement Numbers W911NF-23-2-0216 and W911NF-23-2-0223. The Innsbruck team further received support from the IQI GmbH. M.R. acknowledges support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy Cluster of Excellence Matter and Light for Quantum Computing (ML4Q) EXC 2004/1 390534769.This research is also part of the Munich Quantum Valley (K-8), which is supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of IARPA, the Army Research Office, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. Views and opinions expressed are those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them.
@@
Data and analysis scripts supporting the findings of this study are openly available [46].